Calculate the number of grams of solute necessary to prepare a 500.0 mL of a 0.10M

sodium hydroxide solution.


1
Expert's answer
2022-04-05T08:25:05-0400

CM = n / V

n (NaOH) = CM / V = 0.10 / 0.5000 = 0.2 mol

n = m / M

M (NaOH) = 40 g/mol

m (NaOH) = n x M = 0.2 x 40 = 8 g